### v4

@@ Abstract
Recovery rate prediction plays a pivotal role in bond investment strategies, enhancing risk assessment, optimizing portfolio allocation, improving pricing accuracy, and supporting effective credit risk management. However, forecasting faces challenges like high-dimensional features, small sample sizes, and overfitting. We propose a hybrid Quantum Machine Learning model incorporating Parameterized Quantum Circuits (PQC) within a neural network framework. PQCs inherently preserve unitarity, avoiding computationally costly orthogonality constraints, while amplitude encoding enables exponential data compression, reducing qubit requirements logarithmically. Applied to a global dataset of 1,725 observations (1996–2023), our method achieved superior accuracy (RMSE 0.228) compared to classical neural networks (0.246) and quantum models with angle encoding (0.242), with efficient computation times. This work highlights the potential of hybrid quantum-classical architectures in advancing recovery rate forecasting.
@@ 1 Introduction
Recovery rate prediction is a pivotal element of credit risk management, complementing other key metrics such as Exposure at Default (EAD) and Probability of Default (PD) [7]. While EAD and PD assess the likelihood and extent of credit losses, recovery rates uniquely quantify the proportion of funds recoverable following a default. This metric is particularly valuable for risk assessment, portfolio optimization, and pricing strategies. Despite its critical role, recovery rate prediction has not received commensurate attention in practice. A common approach is to assume a constant recovery rate, typically around 40%, even though empirical data often exhibit a wide range from 0% to 100%, frequently bimodal near 10% and 100% [33, 6, 8, 14]. Such oversimplifications can lead to inaccurate risk evaluations, suboptimal investment decisions, and flawed pricing models, especially in distressed or lower-rated bonds. This lack of focus partly stems from the technical challenges associated with accurate recovery rate modeling.
The technical challenges in forecasting recovery rates are substantial. High-dimensional feature spaces combined with limited datasets often lead to overfitting, a common pitfall in predictive modeling. Classical machine learning approaches have attempted to mitigate this issue through techniques such as Orthogonal Neural Networks (OrthNNs) and Unitary Neural Networks (UNNs) [42, 21, 16, 23, 25], which constrain weight matrices to orthogonal or unitary forms. These methods have demonstrated improved generalization and stability in training. However, maintaining orthogonality during gradient-based training is computationally expensive, often requiring additional steps to re-orthogonalize weights, with time complexity scaling as $O(N^{3})$ for input size $N$. Such limitations underscore the need for novel approaches that can effectively address these challenges.
Quantum Machine Learning (QML) presents a promising alternative to addressing complex computational challenges by integrating Parameterized Quantum Circuits (PQC) as quantum nodes within neural network frameworks. PQCs, which are quantum circuits with tunable parameters, inherently preserve unitarity due to the fundamental principles of quantum mechanics. This intrinsic property eliminates the need for computationally intensive constraints required in classical neural networks to maintain orthogonality or unitarity. By embedding PQCs into neural networks, the resulting models benefit from enhanced generalization and stability, particularly for high-dimensional, small-sample-size datasets prone to overfitting [39]. This unitarity-preserving feature not only simplifies model training but also reduces computational burdens, positioning PQCs as a highly efficient and effective alternative to traditional neural network layers—offering significant potential for advancing machine learning applications.
Quantum Neural Networks (QNNs) have become a prominent tool with applications in quantitative finance among QML approaches. For instance, QNNs have been applied to portfolio optimization, where Quantum Circuit Born Machines outperform classical Restricted Boltzmann Machines [5], and to market forecasting, where Quantum Elman Neural Networks have proven effective for sequential data tasks [22]. Additionally, hybrid QNN models have shown advantages in time series forecasting when implemented on Quantum Processing Units (QPUs) [13, 36]. The Quantum Amplitude Estimation (QAE) algorithm [9], known for its quadratic speedup over classical Monte Carlo techniques [26], presents a promising alternative for probabilistic modeling tasks, such as those encountered in option pricing. Building on the foundations of QAE, other QML approaches, including Quantum Generative Adversarial Networks (qGANs) [44], have gained traction for probabilistic modeling, demonstrating their effectiveness in applications like option pricing. Furthermore, [32] introduced Variational Quantum Amplitude Estimation (VQAE), which combines classical variational optimization with QAE. This approach ensures that the circuit depth remains below a desired threshold, highlighting its potential for practical applications in financial pricing tasks. In fraud detection, QNNs have also demonstrated superior performance, achieving better precision and lower false-positive rates compared to classical methods [20, 43]. Our choice of QNNs for recovery rate forecasting is motivated by their ability to model complex relationships in high-dimensional datasets and their proven versatility across various financial applications.
@@
A critical aspect of QNN is the encoding of classical data into quantum states. A recent study by [38] utilized a QNN integrated with a classical neural network for credit scoring of Small and Medium Enterprises (SMEs), achieving comparable performance to classical models with fewer training epochs. The model employed Angle Data Encoding, which simplifies data preparation and enables shallow circuits suitable for Noise Intermediate Scale Quantum (NISQ) hardware but scales linearly with the feature space, limiting its efficiency for high-dimensional datasets. Angle encoding maps classical data to rotation angles of single-qubit gates [40, 37, 35, 15], requiring a number of qubits proportional to the input size, namely, we need as many qubits as the input size. While this approach simplifies circuit preparation and is feasible for current noisy quantum computers, it lacks scalability. In addition to the Angle Encoding, there is another encoding technique, Amplitude Encoding [40, 10, 37, 35]. The key advantage of the Amplitude Encoding approach is its exponential data compression compared to classical requirements, as the number of required qubits increases only logarithmically with the input size. Specifically, the number of qubits decreases from $N$ to $\log_{2}N$, where $N$ indicates the number of input features. For instance, as illustrated in Figure 1(a) and 1(b), a dataset with four features requires four qubits when using Angle Encoding. In contrast, Amplitude Encoding requires only two qubits to encode the same four features. As the number of features in the dataset increases, the advantage of Amplitude Encoding becomes even more pronounced, enabling efficient scaling for higher-dimensional datasets. Moreover, fewer qubits lead to fewer trainable parameters in the PQC, enhancing computational efficiency.

### v5

@@ Abstract
Recovery rate prediction plays a pivotal role in bond investment strategies by enhancing risk assessment, optimizing portfolio allocation, improving pricing accuracy, and supporting effective credit risk management. However, accurate forecasting remains challenging due to complex nonlinear dependencies, high-dimensional feature spaces, and limited sample sizes—conditions under which classical machine learning models are prone to overfitting. We propose a hybrid Quantum Machine Learning (QML) model with Amplitude Encoding, leveraging the unitarity constraint of Parametrized Quantum Circuits (PQC) and the exponential data compression capability of qubits. We evaluate the model on a global recovery rate dataset comprising 1,725 observations and 256 features from 1996 to 2023. Our hybrid method significantly outperforms both classical neural networks and QML models using Angle Encoding, achieving a lower Root Mean Squared Error (RMSE) of 0.228, compared to 0.246 and 0.242, respectively. It also performs competitively with ensemble tree methods such as XGBoost. While practical implementation challenges remain for Noisy Intermediate-Scale Quantum (NISQ) hardware, our quantum simulation and preliminary results on noisy simulators demonstrate the promise of hybrid quantum-classical architectures in enhancing the accuracy and robustness of recovery rate forecasting. These findings illustrate the potential of quantum machine learning in shaping the future of credit risk prediction.
@@ 1 Introduction
Recovery rate prediction is an essential component of credit risk management, considered alongside key metrics such as Exposure at Default (EAD) and Probability of Default (PD) [10]. While EAD and PD assess the likelihood and magnitude of credit losses, recovery rates measure the proportion of funds that can be recovered following a default. Recovery rates are particularly important for risk assessment, portfolio optimization, and pricing strategies. Despite their importance, recovery rates have received relatively little attention in practice. A common simplifying assumption is to use a constant recovery rate, typically around 40%, despite empirical evidence showing a wide range from 0% to 100%, often with bimodality around 10% and 100% [77, 7, 13, 37]. This oversimplification can result in inaccurate risk assessments, suboptimal investment decisions, and flawed pricing models—particularly for distressed or lower-rated bonds. This limited attention is partly due to the technical challenges involved in accurately modeling recovery rates, including data sparsity, high-dimensional feature spaces, and complex nonlinear dependencies.
There are technical challenges in forecasting recovery rates. Traditional statistical models—such as linear regression, logistic regression, and generalized linear models—have shown limited effectiveness in recovery rate prediction due to their inability to model the highly nonlinear and interactive effects among firm characteristics, macroeconomic indicators, and debt contract features. As highlighted by [1, 79, 70], these models are particularly sensitive to outliers, which are prevalent in distressed debt data, and tend to perform poorly when the underlying data distributions are skewed or exhibit heavy tails—common features in recovery rate datasets. Moreover, their reliance on fixed functional forms and strong distributional assumptions (e.g., normality, homoscedasticity) limits their flexibility in capturing heterogeneity across different sectors or credit environments. These models also scale poorly in high-dimensional settings, where the number of explanatory variables is large relative to the sample size, making them prone to overfitting and unstable parameter estimates.
Classical machine learning approaches have addressed training stability and generalization issues by constraining weight matrices to orthogonal or unitary forms, using techniques such as Orthogonal Neural Networks (OrthNNs) and Unitary Neural Networks (UNNs) [9, 56, 44, 12, 55, 60, 96, 62]. These constraints ensure that each layer acts as a local isometry due to the bounded spectral norm of orthogonal and unitary matrices. [60] showed that optimal generalization is achieved when weight matrices have equal singular values. This reduces sensitivity to perturbations and serves as an implicit regularizer [14, 60]. Theoretically, bounding the spectral norm also lowers capacity measures such as Rademacher complexity and fat-shattering dimension [9, 14]. Empirically, orthogonality has been shown to improve test performance and robustness across tasks [12, 19, 96]. However, enforcing orthogonality during gradient-based training is computationally demanding, often requiring re-orthogonalization steps with time complexity $O(N^{3})$ for input size $N$ [60]. This underscores the need for novel approaches.
The computational challenges in recovery rate prediction stem from two intertwined factors: the inherently complex relationships between financial variables (e.g., non-linear dependencies between macroeconomic indicators and borrower defaults), the small-sample nature of recovery rate data (e.g., limited historical default records for niche markets) which demands a specialized approach to prevent overfitting. Quantum Neural Networks (QNNs) present a promising alternative by leveraging key properties of quantum systems: the exponential growth of Hilbert space with the number of qubits enables QNNs to represent high-dimensional features with fewer parameters, while the intrinsic unitarity of quantum circuits naturally constrains model capacity and helps reduce the risk of overfitting. For instance, [42] demonstrated that a quantum-enhanced kernel method achieves 98% accuracy on a 100-image MNIST subset, outperforming classical SVM (92%), by utilizing entanglement and superposition to efficiently extract patterns from minimal data. Furthermore, [90] theoretically established that QNNs exhibit slower growth of Rademacher complexity with increasing qubits, suggesting superior generalization in low-data scenarios. These unique advantages make QNNs a compelling alternative for prediction problems that combine data scarcity with high-dimensional feature spaces. Motivated by this, our study investigates the feasibility and performance of QNNs in the context of recovery rate prediction—an application that not only poses significant challenges for classical approaches, but also offers an ideal testbed for evaluating the generalization potential of quantum models under realistic constraints.
@@
A critical aspect of Quantum Neural Networks (QNNs) is the encoding of classical data into quantum states. A recent study by [84], through the combination of a QNN with a classical neural network, built a Quantum Machine Learning (QML) model for credit scoring of Small and Medium Enterprises (SMEs). The proposed QML model achieved performance comparable to classical models with fewer training epochs. Their model employed Angle Encoding, which maps classical inputs to rotation angles of single-qubit gates [88, 83, 80, 39]. While this method simplifies circuit design and enables shallow circuits well-suited to current Noisy Intermediate-Scale Quantum (NISQ) hardware, its scalability is limited because the number of required qubits grows linearly with the input dimensionality. As a result, it becomes inefficient for high-dimensional datasets. In contrast, our study adopts Amplitude Encoding, which encodes data into the amplitudes of a quantum state, allowing for exponentially more compact representations [88, 17, 83, 80]. The required number of qubits grows only logarithmically with the input size. This leads to significantly improved scalability, especially for datasets with large feature spaces. Moreover, using fewer qubits results in fewer trainable parameters within the Parameterized Quantum Circuit (PQC), thereby enhancing computational efficiency.
To sum up, in this paper, we propose a Quantum Machine Learning (QML) model that leverages Amplitude Encoding to directly address key challenges in recovery rate prediction. Unlike prior work that primarily uses Angle Encoding [e.g., 87, 61], our approach efficiently encodes high-dimensional financial data into quantum amplitudes, achieving exponential input compression. By integrating amplitude-encoded QNN with classical neural layers, our model achieves parameter efficiency without compromising expressiveness. This hybrid architecture effectively addresses three central challenges: small-sample learning, computational complexity, and relational nonlinearity—challenges that are difficult to overcome using existing classical or quantum methods alone.
The proposed method demonstrated superior performance using a global dataset of 1,725 observations with 256 features spanning 576 firms from 1996 to 2023. It achieved a Root Mean Square Error (RMSE) of 0.228, lower than the RMSE of 0.246 for classical Neural Networks and 0.242 for quantum models with Angle Encoding (see the Results section below for more details). Additionally, the QML model with Amplitude Encoding has fewer trainable parameters, leading to a faster training time of 0.73 seconds per epoch compared to the 0.81 seconds necessary for the QML with Angle Encoding. The lower qubit requirements and reduced computation time underscore the practical applicability of our method for recovery rate forecasting.
This paper makes three main contributions to the literature on QML and financial risk modeling. First, unlike the widely used Angle Encoding, we propose a QML model that uses Amplitude Encoding, more suitable for high-dimensional data representation with fewer qubits [91, 100]. Our work contributes to ongoing research on how different data encoding strategies impact the scalability and practical applicability of quantum models [8, 53].
Second, we apply the QML to the task of predicting recovery rates—an important topic in credit risk modeling. Previous studies have mostly relied on linear models or classical machine learning techniques [6, 79, 72]. In our empirical analysis, the QML with Amplitude Encoding performs better than both traditional models and QML with Angle Encoding. It delivers higher prediction accuracy, more stable results across different runs, and faster convergence, especially when the training data are limited and the relationship between variables is nonlinear.
Third, our approach offers a general modeling tool for financial problems that involve small datasets, high-dimensional features, and complex variable interactions. These conditions are common in many areas of financial risk management and often limit the effectiveness of standard models [16, 11]. By showing that the QML with Amplitude Encoding can work well in such settings, we contribute a new method for analyzing difficult financial prediction problems where data are scarce and patterns are complex.
The remainder of this paper is organized as follows: Section 2 reviews related literature on recovery rate modeling and Quantum Machine Learning (QML). Section 3 presents the data and its features. Section 4 details the hybrid QML approach and Amplitude Encoding. Section 5 discusses the numerical analysis, and Section 6 concludes with insights and directions for future research.
2 Literature Review
Understanding the determinants of recovery rates given default is critical for credit risk modeling, pricing, and risk management. While a vast literature has emerged on modeling the probability of default (PD) (see, e.g., [24]; [29]), the modeling of recovery rates has received relatively limited attention until recently. Yet, recovery rates play a central role in determining the expected loss and are essential for accurate credit risk assessments and financial regulations.
This literature review organizes the related studies into four major streams: (i) conventional econometric modeling techniques; (ii) advanced machine learning methods; (iii) methodological challenges and innovations in high-dimensional recovery modeling and (iv) Quantum Machine Learning applications. We also highlight key differences across these approaches and identify how recent developments contribute to addressing gaps in the literature.
2.1 Traditional Approaches to Recovery Rate Modeling
Traditional recovery rate modeling has primarily relied on econometric techniques that assume linear or parametric relationships between recovery outcomes and explanatory variables. These approaches include linear regression, Tobit models, ordered choice models, and structural credit risk models, each with specific assumptions and limitations.
Linear regression models represent the earliest and most straightforward approach. [6] apply this method using bond data, modeling recovery rates as a function of debt seniority, collateral, and macroeconomic conditions. [1] similarly employ regression analysis to study how governance structures and creditor rights affect recoveries. These models are easy to implement and interpret but fail to address the bounded nature of recovery rates and ignore potential non-linearities.
To address boundary issues, some studies adopt Tobit models. Since recovery rates are censored between 0 and 1, Tobit models are suitable for dealing with observations that pile up at these extremes. [11] offer a comprehensive review of LGD models, advocating for Tobit-based approaches when dealing with censored recovery data. [79] use Tobit specifications to assess the impact of loan and borrower characteristics on recovery outcomes in U.S. bank loans.
Another common method is the use of ordered logit or probit models, which discretize recovery rates into categories. [22] models recovery as an ordered outcome to better handle its bimodal distribution. [23] adopt similar techniques when analyzing recovery rates of European bank loans, emphasizing loan type and provisioning practices.
Beyond these reduced-form approaches, structural credit risk models offer a more theoretical perspective. These models derive recovery as an endogenous outcome of firm value at the time of default. [30] develop a structural model in which recovery depends on firm leverage and the timing of default. [38] extend the Merton framework by integrating dynamic firm fundamentals and market conditions. [27] implement a structural framework that emphasizes the importance of credit cycle effects in explaining recovery rate dynamics.
These approaches often suffer from limitations in flexibility and scalability. The parametric structure requires pre-specification of functional forms, and the small number of predictors leaves out potentially important nonlinear effects and interaction terms. This has led to a growing interest in semi-parametric or non-parametric methods that can better accommodate the complex structure of the recovery process.
However, these methods still face significant challenges when dealing with high-dimensional predictor spaces. Additionally, they often require strong parametric assumptions and may not generalize well across datasets with small sample sizes—a persistent issue in recovery rate research due to the rarity of default events.
In summary, traditional models offer important insights into recovery processes but suffer from rigidity in functional form, limited capacity to capture complex interactions, and challenges in accommodating high-dimensional data. These limitations motivate the use of more flexible and scalable machine learning methods, as explored in this project.
2.2 Machine Learning Methods in Recovery Modeling
Recent years have witnessed growing interest in Machine Learning (ML) methods for recovery modeling, driven by the increasing availability of high-dimensional and unstructured data, as well as the limitations of traditional approaches. ML models are particularly well-suited to capture nonlinear relationships, complex interactions, and heterogeneity in the data without requiring pre-specified functional forms.
Support Vector Machines (SVMs) and k-Nearest Neighbor (k-NN) classifiers have also been applied in the credit risk domain, although they are less common in recovery modeling specifically. ML models often outperform traditional approaches in terms of predictive accuracy but face challenges related to overfitting, interpretability, and the curse of dimensionality.
Some recent work explores text-based features, such as sentiment extracted from financial news or filings, to improve prediction performance. For instance, [5] and [69, 68] apply NLP techniques in credit scoring, and similar ideas are adapted in this project to model recovery using risk sentiment indices derived from news.
Tree-based ensemble methods such as random forests and gradient boosted trees (e.g., XGBoost) have been used to model recovery rates due to their ability to handle variable interactions and nonlinearity [45, 71, 58].
Neural networks and deep learning techniques [35] offer even greater modeling flexibility, although their application is limited by data sparsity in credit markets. In practice, the small sample size of defaults and limited labeled data on recovery rates make the application of deep learning challenging in credit markets.
We provide an overview of key studies modeling recovery rates for U.S. corporate bonds in Table 1. Existing literature has adopted a wide range of methodologies, which have evolved significantly over time.
Study Methodology Model Type (Key Technique)
[2] Linear Regression Traditional econometric
[92] LASSO Regularization
[41] Linear Regression Traditional econometric
[33] Conditional model Traditional econometric
[34] Probit / Macro analysis Traditional
[94] Univariate and multivariate regression Traditional
[99] Group LASSO Structured Regularization
[1] Regression Traditional econometric
[45] Random Forests / Boosting ML (tree-based)
[79] Quantile Regression Semi-parametric
[47] Beta-link generalized linear model Semi-parametric
[11] Tobit Traditional
[32] Regression with market variables Traditional econometric
[48] Zero-One Inflated Beta Semi-parametric
[22] Ordered Logit Traditional
[27] Structural Credit Risk Model Traditional with macro factors
[69] Text Mining + ML Machine Learning (NLP)
[3] Parametric regressions, regression trees Ensemble + Traditional
[25] Univariate and multivariate regression Traditional
[65] Univariate and multivariate regression Traditional
[16] Best Subset Selection Regularization / Optimization
[35] Neural Networks Deep Learning
[52] Multivariate linear regression Traditional
[57] SHAP values Model interpretation
[50] Inverse Gaussian, regression trees, mixtures Semi-parametric / ML
[70] SVRs, bagging, boosting, LASSO, ridge Machine Learning
[5] NLP-based Sentiment Machine Learning (BERT)
[36] Beta regression, mixture models, regression trees Semi-parametric + ML
Table 1: Summary of Recovery Rate Modeling Approaches in the Literature.
2.3 Quantum Machine Learning
Quantum Machine Learning (QML) models combine Quantum Neural Networks (QNNs) with classical layers. This new emerging field, leveraging the power of quantum computing with Machine Learning techniques, addresses problems commonly challenging for classical algorithms. Thanks to quantum parallelism, QML models have attracted attention for their potential to process high-dimensional data and capture complex nonlinearities in a more compact architecture. QNNs leverage quantum properties, including superposition, entanglement, and unitary transformations, to encode and transform input data via a Parameterized Quantum Circuit (PQC). These models are especially promising in scenarios where classical models suffer from the curse of dimensionality, such as in high-dimensional recovery rate modeling with sparse default data.
Recent studies have explored the application of QML in finance and credit risk. For example, [91] and [18] provide foundational discussions on how quantum algorithms can be adapted to supervised and unsupervised learning tasks. More recent work by [8] and [17] explores how QNNs can approximate functions in complex spaces and perform optimization tasks over rugged objective surfaces. These features make QNNs especially suitable for credit risk modeling, where the signal-to-noise ratio is low and default-related data is sparse. In practice, quantum-enhanced models such as Variational Quantum Circuits (VQCs), Quantum Boltzmann Machines (QBM), and Quantum Convolutional Neural Networks (QCNNs) have shown success in tasks like classification and regression with limited data [4] and success in market forecasting, where Quantum Elman Neural Networks have proven effective for sequential data tasks [59]. Similarly, [42] shows that quantum models can perform competitively on structured prediction tasks and kernel-based learning in domains with limited training observations. Additionally, QML models have shown advantages in time series forecasting when implemented on Quantum Processing Units (QPUs) [31, 82]. The Quantum Amplitude Estimation (QAE) algorithm [15], known for its quadratic speedup over classical Monte Carlo techniques [63], presents a promising alternative for probabilistic modeling tasks, such as those encountered in option pricing. Building on the foundations of QAE, other QML approaches, including Quantum Generative Adversarial Networks (qGANs) [100], have gained traction for probabilistic modeling, demonstrating their effectiveness in applications like option pricing. Furthermore, [75] introduced Variational Quantum Amplitude Estimation (VQAE), which combines classical variational optimization with QAE. This approach ensures that the circuit depth remains below a desired threshold, highlighting its potential for practical applications in financial pricing tasks. In fraud detection, QNNs have also demonstrated superior performance, achieving better precision and lower false-positive rates compared to classical methods [53, 93].
As quantum hardware continues to evolve, QNNs offer a pathway to overcoming data limitations and complexity in financial modeling. In this paper, we adopt a QML model to capture complex nonlinear interactions in high-dimensional feature spaces and to enhance the recovery rate prediction under data-scarce conditions. By leveraging hybrid quantum-classical architectures, our approach aims to exploit the representational power of QML while maintaining computational tractability on near-term quantum devices.
3 Data
We consider a dataset comprising 256 features and 1,725 observations covering 576 firms from 1996 to 2023. The data is obtained through the NRF Research Project UP5 of the National University of Singapore. The UP5 data contains Macroeconomic and market-related features obtained from FRED and Refinitiv; financial statement features of firms sourced from Bloomberg; and bond-level features, as well as firm-level or market-level credit product features provided by the Credit Research Initiative (CRI) of the National University of Singapore.
3.1 Summary Description of Recovery Rate
[…]
Obs. Mean STD Min 25% 50% 75% Max
Recovery Rate 1,725 0.4845 0.3317 0 0.1811 0.4170 0.7896 1.0996
Table 2: Summary Statistics of Recovery Rate
3.2 Overview of Features
In total, we construct 256 features that encompass firm-level characteristics, bond-specific attributes, and macroeconomic indicators. These features are selected based on their relevance to credit risk and recovery rate modeling, as supported by prior literature such as [49] and [40]. For ease of presentation and interpretation, we organize these features into several broad categories as shown in Table 3.
Category Examples Count
Bond-Specific Variables Coupon rate, duration, seniority, default event type 29
Firm-level Characteristics Return on assets (ROA), debt-to-equity ratio, current ratio 17
Credit Risk Measures Probability of default (PD) in 1 month, distance to default (DTD), sigma 78
Industry, Geographic, and Currency Variables Industry dummies, region identifiers, currency indicators 98
Macroeconomic and Market Conditions GDP growth, inflation, financial uncertainty indices, treasury yields 34
Table 3: Summary of feature categories used in recovery rate prediction. The table groups the 256 features into five major categories and provides illustrative examples for each type.
The first group contains bond-specific variables. These include information that is intrinsic to the bond contract itself, such as the coupon rate, duration, and seniority level. We also include a series of event indicators capturing credit events like defaults, restructurings, delistings, and insolvencies. In total, this category comprises 29 variables. The data for this group are primarily obtained from Bloomberg and Refinitiv.
The second group focuses on firm-level financial indicators. These variables are constructed using company financial statements and include standard measures such as return on assets (ROA), debt-to-equity ratio, current ratio, and market capitalization. These metrics provide insight into the issuing firm’s solvency and operating health. This category includes 17 variables and is sourced entirely from Bloomberg.
The third category consists of credit risk measures, particularly probability of default (PD) estimates. We incorporate a comprehensive range of PD scores calculated across different aggregation levels, including sector, domicile, exchange, and global scopes. These variables are forward-looking and help capture variation in credit quality across firms. In total, we include 77 PD-related variables, all obtained from the Credit Research Initiative (CRI) of the National University of Singapore.
The fourth group includes categorical variables that account for structural heterogeneity across firms. Specifically, we introduce one-hot encoded dummies for industries, regions, and currencies. These variables allow us to control for country-specific and sector-specific effects. There are 55 industry dummies and 43 regional or currency indicators, summing to 98 variables in this category. These are also provided by the CRI.
Finally, we include a group of macroeconomic and market-related variables. This group captures the broader economic environment and includes variables such as interest rates of various maturities, GDP growth, inflation, financial uncertainty indices, and measures of market volatility. These 34 variables are collected from FRED and Bloomberg, and they serve to contextualize firm and bond behavior within the prevailing macro-financial landscape.
All features are selected based on domain expertise and prior empirical findings that link these attributes to credit outcomes and recovery dynamics.
4 Methodology
Our proposed Quantum Machine Learning (QML) model adopts a hybrid architecture, consisting of a classical preprocessing layer followed by a Quantum Neural Network (QNN) (the overall architecture is depicted in Figure 4). This design is inspired by [84], who demonstrated the effectiveness of integrating classical neural components with quantum circuits in financial classification task. The overall pipeline of our hybrid QML model comprises the following four steps:
  1. 1.
    Classical Preprocessing (pink block in Figure 4): The raw input features ($N=2^{n}$ dimensions) are first processed by a classical neural network, which acts as a feature transformation or embedding layer to condition the data for quantum encoding.
  2. 2.
    Quantum State Preparation (connection arrow between the pink block and the dashed block inside the purple block in Figure 4, and Figure 5(b)): The transformed features are encoded into a quantum state |$\ket{\psi}$ via Amplitude Encoding. This encoding scheme efficiently compresses high-dimensional data, requiring only $n=\log_{2}N$ qubits for $N$ features.
  3. 3.
    Parametrized Quantum Circuit (the solid block inside the purple block in Figure 4): The encoded quantum state is then passed through a Parameterized Quantum Circuit (PQC), composed of rotation gates and CNOT gates which produce entanglement.
  4. 4.
    Measurement and Output: After quantum processing, the qubits are measured in the computational basis. The measurement results are forwarded to a classical output layer to generate the final prediction (i.e., the recovery rate).
While the execution of QML models on real quantum hardware remains constrained by noise, decoherence, and limited qubit counts, this study is conducted using a fault-tolerant quantum simulator. This allows us to isolate and evaluate the algorithmic performance of the proposed architecture without the confounding effects of hardware-induced errors.
[…]
QNN Amplitude EncodingOutput
4.1 Classical Preprocessing
As shown in Figure 4, the input classical data—comprising $N=2^{n}$ features—is first processed by a classical neural network (pink block). The classical master layer comprises an input layer and a hidden layer of equal size, using a LeakyReLU activation function. This architecture offers three advantages. First, it extracts meaningful internal representations from high-dimensional, redundant input data. Second, it acts as a feature transformation or embedding layer to prepare the data for quantum data encoding. Third, it introduces non-linearity, enabling the model to capture complex relationships.
4.2 Quantum Data Encoding
To interface between the classical and quantum layers, we encode the transformed classical data into a quantum state, denoted as |$\ket{\psi}$ (see Figures 4 and 5(b)).
Two commonly used methods for quantum data encoding are Angle Encoding and Amplitude Encoding, illustrated in Figures 5(a) and 5(b), respectively. In Angle Encoding, classical feature values are mapped to the rotation angles of single-qubit gates (Figure 5(a)). While straightforward, this method requires one qubit per input feature, making it impractical for high-dimensional datasets such as those used in recovery rate prediction.
To overcome this limitation, we adopt Amplitude Encoding, which maps a $2^{n}$-dimensional classical feature vector into the amplitude components of an $n$-qubit quantum state as follows
|$\ket{\psi}=\sum_{i=0}^{N-1}\alpha_{i}\ket{b_{1}^{i}\dots b_{n}^{i}},$ (1)
where $N=2^{n}$ is the number of features, $b_{1}^{i},\dots,b_{n}^{i}$ is the binary representation of integer $i$, and $\alpha_{i}$ are the normalized amplitudes derived from the input data.